A bus leaves a station at 1 PM,traveling west at an average rate of 44mi/hour. One hour later a second bus leaves the station, traveling east at a rate of 48 mi/h. At what time will the two buses be 274 miles apart?

Answer by stanbon(75887)

A bus leaves a station at 1 PM,traveling west at an average rate of 44mi/hour. One hour later a second bus leaves the station, traveling east at a rate of 48 mi/h. At what time will the two buses be 274 miles apart?
------------------------------
West-bound DATA:
Rate= 44 mph ; Time = x hrs. ; distance = 44x miles
-------------------------------------
East-bound DATA:
Rate = 48 ph ; Time = x-1 hrs ; distance = 48(x-1) miles
------------------------------------
EQUATION:
distance + distance = 274
44x + 48(x-1) = 274
44x + 48x - 48 = 274
92x = 322
x = 3.5 hrs
--------------------
1PM = 3.5 hrs = 4:30PM
========================================
Cheers,
Stan H.